# Demonstration of weighted graph optimization on a Rydberg atom array using local light-shifts

### v1

@@ Abstract
Neutral atom arrays have emerged as a versatile platform towards scalable quantum computation and optimization. In this paper we present first demonstrations of weighted graph optimization on a Rydberg atom array using annealing with local light-shifts. We verify the ability to prepare weighted graphs in 1D and 2D arrays, including embedding a five vertex non-unit disk graph using nine physical qubits. We find common annealing ramps leading to preparation of the target ground state robustly over a substantial range of different graph weightings. This work provides a route to exploring large-scale optimization of non-planar weighted graphs relevant for solving relevant real-world problems.
@@
The ability to implement programmable spin models extends the utility of neutral atom arrays to the regime of analogue quantum computing [33], where in the context of graph optimization coupling atoms to the Rydberg states enables native embedding of the Maximum Independent Set (MIS) problem for planar 2D unit-disk graphs (UDG) [34, 35]. Here the Rydberg blockade mechanism [36] prevents more than a single Rydberg excitation occuring within a radius $r<r_{\mathrm{B}}$, enforcing the constraint that no two connected vertices should be simultaneously part of the same MIS configuration. Experimental demonstrations of solving MIS on 2D-UDG using QAA with hundreds of qubits demonstrate the feasibility of this approach [37, 38], however recent analysis shows for the native Kings-graph classical solvers are able to find solutions to graphs with 1000 vertices in minutes [39]. One approach for extending beyond planar 2D graph is to use ancillary qubits to create quantum wires [40] which can be routed through neighboring trapping planes in 3D to implement non-local graph connectivity [41]. This approach has been used to demonstrate embedded integer factorization [42] and quadratic unconstrained binary optimization (QUBO) problems [43] onto atom arrays using only global control.
@@
Another approach to extend the range of problems that can be encoded using neutral atom arrays is to introduce local control fields. This permits the introduction of weighting to enable solution of Maximum Weighted Independent Set (MWIS) problems on a 2D UDG, providing a route to map a variety of graph and QUBO problems onto neutral atom arrays using this MWIS-UDG encoding [44] with at worst a quadratic overhead in physical qubit number. Similar scaling is obtained using a related parity-based encoding for graph problems [45], whilst moving beyond the blockade regime to remove the UDG restriction offers prospects for linear scalings in qubit overhead [46]. To date, local light-shifts have been used on neutral atom arrays in the field of quantum simulation for generating large-scale entangled states [47] and implementing fast, arbitrary local control of interacting trimers [48].
@@
Experimental Setup.— The experimental setup is based on arrays of holographically trapped Cs atoms. Atoms are initially prepared in the $|g\rangle=|4,0\rangle$ state, and are globally excited to the |$\ket{r}=\ket{80s_{1/2},m_{j}=+1/2}$ Rydberg state via two-photon excitation using the $7p_{1/2}$ intermediate state using light at 459 nm and 1039 nm. We operate with a detuning of $\Delta^{\prime}/2\pi=$500 MHz from the intermediate state, and achieve a Rydberg Rabi frequency of $\Omega/2\pi=2.70(2)$ MHz. Global detuning is controlled by adjusting the frequency using a double-passed acousto-optic modulator (AOM) in the 459 nm path. For $80s_{1/2}$ we calculate $C_{6}=-3376~{}\mathrm{GHz}\,\mu$m6 [51], resulting in a resonant blockade radius $r_{B}=\sqrt[6]{|C_{6}|/\Omega}=10.4~{}\mu$m.
Local light-shifts are realized using light at 800 nm derived from a Ti:Sapph laser that is projected from a spatial light modulator (SLM) and overlapped with the main optical trapping array. To suppress mechanical forces coming from the repulsive optical potential, the light-shift spot arrays are re-imaged onto the atoms with a $1/e^{2}$ radius of $5.1(2)~{}\mu$m. For a given UDG-MWIS, the arrays of local light-shift beams are generated using a Gerchberg-Saxon algorithm with the desired relative weighting [52]. To verify the light-shifts experienced by the atoms, we perform spectroscopy on the |$\ket{g}\rightarrow\ket{50s_{1/2},m_{j}=+1/2}$ transition to suppress interactions, and perform a closed-loop feedback on the SLM potentials to achieve target weightings with $<2\%$ RMS errors in the relative weightings after 5 iterations  [50]. The resulting differential light-shift is calibrated as $\delta_{\mathrm{AC}}=0.28(2)$ MHz///mW, which is dominated by the ground state light-shift as the Rydberg AC shift calculated using the polarizability of a free electron corresponds to only 40 kHz [53].
@@
Discussion and Outlook In this paper we have demonstrated a new approach to implementing UDG-MWIS using arbitrary local light-shifts implemented using a SLM combined with a dual stage annealing profile that combines the advantages of local light-shifts with global frequency and intensity control for implementing a quantum annealing algorithm. We demonstrate the ability to prepare ground states of small-scale 1D and 2D graphs, including demonstrating the embedding of a non-UDG MWIS problem. In both cases, we were able to identify robust annealing protocols suitable for use with different relative weightings providing the possibility to design universal ramps that are suitable for solving a large number of graphs without the required overhead of full closed-loop optimization of the annealing ramp each time.
@@ Acknowledgements.
We thank H. Pichler for helpful discussions. This work is supported by the EPSRC Prosperity Partnership SQuAre (Grant No. EP/T005386/1) with funding from M Squared Lasers Ltd. The data presented in this work are available at [56].
@@
Figure 5: Experimental Setup. (a) Atom arrays are generated using 1064 nm trapping light projected into the chamber using an SLM and a high-numerical aperture in-vacuum aspheric lens. Local light-shifts are implemented using a second SLM to overlap spots at 800 nm with the 1064 nm tweezers. Atom imaging is performed by collecting fluoresence onto an sCMOS camera. Rydberg excitation is performed using counter-propagating beams at 459 nm and 1039 nm with orthogonal circular polarisiations which globally illuminate the array. (b) Two-photon Rydberg excitation scheme from the $6S_{1/2}$ clock state $|g\rangle=|4,0\rangle$ to $|r\rangle=|ns_{1/2},m_{j}=+1/2\rangle$ via the intermediate $7P_{1/2}$ level. Lasers are locked with a detuning of $\Delta^{\prime}/2\pi=500$ MHz from the $7P_{1/2}$ transition.
@@ I.2 Rydberg Excitation
The 459 nm light is frequency doubled (M Squared ECDF) and the detuning controlled using a double pass accousto-optic modulator (AOM,AA MQ180-A0.25-VIS) in a cat-eye geometry to enable fast frequency control. The AOM is driven using a high bandwidth arbitrary waveform generator (AWG, Spectrum M4i6631-x8). This is delivered to the experiment by fiber, and overlapped onto the array with a 125 $\mu$m $1/e^{2}$ waist. The 1039 nm laser seeds a fiber amplifier (Azur ALS-1039-20-CP) which is intensity controlled using a single-pass AOM and coupled into a high-power fiber. The beam is then imaged onto the atoms using cylindrical optics to create a beam with a $1/e^{2}$ waist size of $60\times 20~{}\mu$m. For the experiments presented here, Rydberg excitation is performed using an intermediate detuning of $\Delta^{\prime}/2\pi=500$ MHz with respect to the $7P_{1/2}$ centre of mass energy, with powers set to provide a Rabi frequency of $\Omega/2\pi=2.70(0)$ MHz over the central rows of the array.
@@ I.3 Local Light-shifts
For performing local-light shifts, we suppress the mechanical effects of the repulsive light shift on the ground-state atoms by changing the input beam size on the SLM to create spots with an effective waist size of 5.1 $\mu$m, resulting in $<2\%$ additional atom loss when applying local shifts of up to 10 MHz for periods of up to 3 $\mu$s.

### v2

@@ Abstract
Neutral atom arrays have emerged as a versatile platform towards scalable quantum computation and optimization. In this paper we present first demonstrations of weighted graph optimization on a Rydberg atom array using annealing with local light-shifts. We verify the ability to prepare weighted graphs in 1D and 2D arrays, including embedding a five vertex non-unit disk graph using nine physical qubits and demonstration of a simple crossing gadget. We find common annealing ramps leading to preparation of the target ground state robustly over a substantial range of different graph weightings. This work provides a route to exploring large-scale optimization of non-planar weighted graphs relevant for solving relevant real-world problems.
@@
The ability to implement programmable spin models extends the utility of neutral atom arrays to the regime of analogue quantum computing [33, 34], where in the context of graph optimization coupling atoms to the Rydberg states enables native embedding of the Maximum Independent Set (MIS) problem for planar 2D unit-disk graphs (UDG) [35, 36]. Here the Rydberg blockade mechanism [37] prevents more than a single Rydberg excitation occuring within a radius $r<r_{\mathrm{B}}$, enforcing the constraint that no two connected vertices should be simultaneously part of the same MIS configuration. Experimental demonstrations of solving MIS on 2D-UDG using QAA with hundreds of qubits demonstrate the feasibility of this approach [38, 39], however recent analysis shows for the native Kings-graph classical solvers are able to find solutions to graphs with 1000 vertices in minutes [40]. One approach for extending beyond planar 2D graph is to use ancillary qubits to create quantum wires [41] which can be routed through neighboring trapping planes in 3D to implement non-local graph connectivity [42]. This approach has been used to demonstrate embedded integer factorization [43] and quadratic unconstrained binary optimization (QUBO) problems [44] onto atom arrays using only global control.
@@
Another approach to extend the range of problems that can be encoded using neutral atom arrays is to introduce local control fields as initially proposed in [45]. This permits the introduction of weighting to enable solution of Maximum Weighted Independent Set (MWIS) problems on a 2D UDG, providing a route to map a variety of graph and QUBO problems onto neutral atom arrays using this MWIS-UDG encoding [46] with at worst a quadratic overhead in physical qubit number. Similar scaling is obtained using a related parity-based encoding for graph problems [47]. Moving beyond the blockade regime to remove the UDG restriction offers prospects for a linear scaling in qubit overhead and application to MaxCut problems [45]. To date, local light-shifts have been used on neutral atom arrays in the field of quantum simulation for generating large-scale entangled states [48], implementing fast, arbitrary local control of interacting trimers [49] and studying quantum coarsening dynamics [50].
@@
Experimental Setup.— The experimental setup is based on arrays of holographically trapped Cs atoms. Atoms are initially prepared in the $|g\rangle=|4,0\rangle$ state, and are globally excited to the |$\ket{r}=\ket{80s_{1/2},m_{j}=+1/2}$ Rydberg state via two-photon excitation using the $7p_{1/2}$ intermediate state using light at 459 nm and 1039 nm. We operate with a detuning of $\Delta^{\prime}/2\pi=$502 MHz from the intermediate state, and achieve a Rydberg Rabi frequency of $\Omega/2\pi=2.70(2)$ MHz. Global detuning is controlled by adjusting the frequency using a double-passed acousto-optic modulator (AOM) in the 459 nm path. For $80s_{1/2}$ we calculate $C_{6}=-3376~{}\mathrm{GHz}\,\mu$m6 [53], resulting in a resonant blockade radius $r_{B}=\sqrt[6]{|C_{6}|/\Omega}=10.4~{}\mu$m.
Local light-shifts are realized using light at 800 nm derived from a Ti:Sapph laser that is projected from a spatial light modulator (SLM) and overlapped with the main optical trapping array. To suppress mechanical forces coming from the repulsive optical potential, the light-shift spot arrays are re-imaged onto the atoms with a $1/e^{2}$ radius of $3.0(4)~{}\mu$m. For a given UDG-MWIS, the arrays of local light-shift beams are generated using a Gerchberg-Saxon algorithm with the desired relative weighting [54]. To verify the light-shifts experienced by the atoms, we perform spectroscopy on the |$\ket{g}\rightarrow\ket{50s_{1/2},m_{j}=+1/2}$ transition to suppress interactions, and perform a closed-loop feedback on the SLM potentials to achieve target weightings with $<2\%$ RMS errors in the relative weightings after 5 iterations  [52]. The resulting differential light-shift is calibrated as $\delta_{\mathrm{AC}}=0.28(2)$ MHz///mW, which is dominated by the ground state light-shift as the Rydberg AC shift calculated using the polarizability of a free electron corresponds to only 40 kHz [55].
@@
Gadgets.— As discussed above, a number of proposals exist to extend the range of target optimisation problems that can be solved using neutral atom arrays by converting to a UDG-MWIS [45, 46, 47]. One such approach introduces simple gadgets that provide a generalised framework for mapping problems to UDG-MWIS [46], and here we consider a proof-of-concept implementation of the crossing gadget using our local light-shift approach. This crossing gadget, shown in Fig. 5(a), implements an edge between qubits 1 and 2, and qubits 3 and 4. This requires four additional atoms, where the logical qubits (1-4) are assigned weight $w_{1}=1$, whilst the ancilla gadgets have $w_{2}=2.86$ to ensure the lowest energy state corresponds to an equal superposition of the four degenerate output configurations with either qubit 1 or 2 excited independent of 3 or 4. This deviates from $w_{2}=4$ presented in [46] due to the requirement to compensate for the effects of the $1/R^{6}$ interaction tails on the ancilla atoms. For these measurements the intermediate state detuning was increased to $\Delta^{\prime}/2\pi=1$ GHz and Rabi frequency reduced to $\Omega/2\pi=1.28$ MHz motivated by numerical modelling of the annealing process. As above, we then optimize the annealing ramp experimentally to minimize $\langle H_{\mathrm{MWIS}}\rangle$ for the problem resulting in the ramp shown in Fig. 5(b). The distribution of the final output strings are shown in Fig. 5(c), where the four target ground states are the most likely to be observed with approximately equal probability, and $\geq 20\%$ combined probability. Analysis of the next four most probable output strings show these correspond to the case of a target ground state with one of the logical qubit Rydberg excitation’s missing, likely due to limitations in the Rydberg detection fidelity. For embedding of larger graphs, improvements in the Rydberg state detection can be implemented using microwave ionization [56, 29], however these results verify the proof-of-concept approach for performing gadget encodings on the current architecture.
Figure 5: Crossing gadget. (a) Crossing gadget designed to enable logical crossing of two quantum wires connecting qubits 1 and 2, and 3 and 4, without introducing coupling between them. To compensate for interaction tails we use $w_{1}=1$ for logical qubits and $w_{2}=2.86$ for the ancilla atoms. The spacing between two adjacent sites with $w_{2}$ is $6~{}\mu m$, and $7.5~{}\mu m$ between them and their nearest $w_{1}$ neighbour. (b) Optimized annealing ramp (c) Final gadget output strings showing the four degenerate ground states are each prepared with probabilities of $5(1)\%$.
Discussion and Outlook.— In this paper we have demonstrated a new approach to implementing UDG-MWIS using arbitrary local light-shifts implemented using a SLM combined with a dual stage annealing profile that combines the advantages of local light-shifts with global frequency and intensity control for implementing a quantum annealing algorithm. We demonstrate the ability to prepare ground states of small-scale 1D and 2D graphs, including demonstrating the embedding of a non-UDG MWIS problem. In both cases, we were able to identify robust annealing protocols suitable for use with different relative weightings providing the possibility to design universal ramps that are suitable for solving a large number of graphs without the required overhead of full closed-loop optimization of the annealing ramp each time.
@@ Acknowledgements.
We thank H. Pichler for helpful discussions. This work is supported by the EPSRC Prosperity Partnership SQuAre (Grant No. EP/T005386/1) with funding from M Squared Lasers Ltd and EPSRC Grant EP/Y005058/2. The data presented in this work are available at [59].
@@
Figure 6: Experimental Setup. (a) Atom arrays are generated using 1064 nm trapping light projected into the chamber using an SLM and a high-numerical aperture in-vacuum aspheric lens. Local light-shifts are implemented using a second SLM to overlap spots at 800 nm with the 1064 nm tweezers. Atom imaging is performed by collecting fluoresence onto an sCMOS camera. Rydberg excitation is performed using counter-propagating beams at 459 nm and 1039 nm with orthogonal circular polarisiations which globally illuminate the array. (b) Two-photon Rydberg excitation scheme from the $6S_{1/2}$ clock state $|g\rangle=|4,0\rangle$ to $|r\rangle=|ns_{1/2},m_{j}=+1/2\rangle$ via the intermediate $7P_{1/2}$ level. Lasers are locked with a detuning of $\Delta^{\prime}/2\pi=506$ MHz from the $7P_{1/2}$ transition.
@@ I.2 Rydberg Excitation
The 459 nm light is frequency doubled (M Squared ECDF) and the detuning controlled using a double pass accousto-optic modulator (AOM,AA MQ180-A0.25-VIS) in a cat-eye geometry to enable fast frequency control. The AOM is driven using a high bandwidth arbitrary waveform generator (AWG, Spectrum M4i6631-x8). This is delivered to the experiment by fiber, and overlapped onto the array with a 125 $\mu$m $1/e^{2}$ waist. The 1039 nm laser seeds a fiber amplifier (Azur ALS-1039-20-CP) which is intensity controlled using a single-pass AOM and coupled into a high-power fiber. The beam is then imaged onto the atoms using cylindrical optics to create a beam with a $1/e^{2}$ waist size of $60\times 20~{}\mu$m. For the experiments presented here, Rydberg excitation is performed using an intermediate detuning of $\Delta^{\prime}/2\pi=506$ MHz with respect to the $7P_{1/2}$ centre of mass energy, with powers set to provide a Rabi frequency of $\Omega/2\pi=2.70(0)$ MHz over the central rows of the array.
@@ I.3 Local Light-shifts
For performing local-light shifts, we suppress the mechanical effects of the repulsive light shift on the ground-state atoms by changing the input beam size on the SLM to create spots with an effective waist size of 3.0(4) $\mu$m, resulting in $<2\%$ additional atom loss when applying local shifts of up to 10 MHz for periods of up to 3 $\mu$s.